alones (alones):

@Agent0smith please don't lurk here :3

alones (alones):

\[5x^2+2x-4=0 \] OG SAVE ME, I NEED THIS CHECKED

rebeccaxhawaii (rebeccaxhawaii):

Thats all he does. He has 9054905483980498 tabs open dedicated to lurk.

alones (alones):

o.m.g damn.

rebeccaxhawaii (rebeccaxhawaii):

I hate mud sticks -.- but we are having fun while doing math. Is it a crime?

rebeccaxhawaii (rebeccaxhawaii):

Anyways what have you got so far?

alones (alones):

My answer is \[\frac{ -1+\sqrt{21} }{ 5 },\frac{ -1-\sqrt{21} }{ }\]

OpenStudy (agent0smith):

Yes.

alones (alones):

THANK YOU

OpenStudy (agent0smith):

You're welcome.